OpenStudy (anonymous):

What are the solutions of x^2 + 5x − 24 = 0?

OpenStudy (anonymous):

Best bet is a graphing calculator. The solutions will be wherever the graph crosses the x-axis. Also, you can factor that equation to find the solutions. Find 2 numbers that will multiply to 24 and subtract to 5. Do you know what to do from there or do you need more help?

hero (hero):

x^2 + 8x - 3x - 24 = 0 x(x+8)-3(x+8)=0 (x+8)(x-3) = 0 x = -8,3